Present Value of $1 at Compound Interest EXHIBIT 8 Periods 4% 4y% 5% 5,496 6% 6Y5% 7% 10% 11% 12% 13% 1 0.96154 0.95694 0.95238 0.94787 0.94340 093897 0.93458 090909 0.90090 0.89286 088496 0.92456 091573 0.90703 0.89845 0.89000 088166 0.87344082645 0.81162 0.797190.78315 0.88900 0.87630 0.86384 0.85161 0.83962 082785 0.81630 0.75131 0.73119 0.71178 0.69305 0.85480 0.83856 0.82270 0.807220.79209 0.77732 0.76290 0.68301 0.658730.63552 0.61332 0.82193 0.80245 0.78353 0.76513 0.74726 0.72988 0.71299 0.62092 059345 0.56743 0.54276 6 0.79031 0.76790 0.74622 0.72525070496 068533。166634 0.56447 0534 64 0.50653 048032 7 0.75992 0.73483 0.71068 068744 0.66506 0.6435 0.62275 0.51316 048166 045235 042506 8 0.73069 0.70319 0.67684 065160 0.62741 060423 0.5820 0.46651 043393 040388 0.37616 9 0.70259 0.67290 64461 0.61763 059190 056735 0.54393 0.42410 0.39092 0.36061 0.33288 2 4 10 0.67556 0.64393 0.61391 058543 0.55839 0.53273 0.50835 0.38554 0.35218 0.32197 029459EXHIBIT 1 Present Value of an Annuity of $1 at Compound Interest Periods 4% 42% 5% 5Y296 6% 10% 6 2 7% 11% 12% 1396 0.96154 0.95694 0.95238 0.94787 0.94340 0.93897 0.93458 0.90909 0.90090 0.89286 0.88496 1.88609 87267 1.85941 84632 183339 182063 80802 .73554 .71252 1.69005 166810 2.77509 2.74896 2.72325 2.69793 2.67301 2.64848 2.62432 248685 244371 240183 2.36115 3.62990 3.58753 3.54595 3.50515 346511 3.42580 3.38721 3.169873.10245 3.03735 2.97447 445182 4.38998 4.32948 4.27028 4.21236 4.15568 4.10020 3.79079 3.69590 3.60478 3.51723 5.24214 5.15787 5.07569 4.99553 491732 4.84101 4.76654 4.35526 4.23054 4.111 3.99755 6.00205 5.89270 5.786375.68297 5.58238 5.48452 5.38929 486842 4.71220 4.56376 442261 6.73274 6.59589 6.463216.33457 6.20979 6.08875 5.97130 5.33493 5.14612 4.96764 4.79677 43533 7.26879 7.10782 6.95220 6.80169 6.65610 6.51523 5.75902 5.53705 5.32825 5.13166 8.11090 7.91272 7.721737.53763 736009 7.18883 7.02358 6.144575.88923 5.65022 542624 2 2 10

Compute Bond Proceeds, Amortizing Premium by Interest Method, and Interest Expense Ware Co. produces and sells motorcycle parts. On the first day of its fiscal year, Ware Co. issued $50,000,000 of four-year, 12% bonds at a market (effective) interest rate of 10%, with interest payable semiannually. Compute the following:

a. The amount of cash proceeds from the sale of the bonds. Use the tables of present values in Exhibit 8 and Exhibit 10. Round to the nearest dollar. $

b. The amount of premium to be amortized for the first semiannual interest payment period, using the interest method. Round to the nearest dollar. $

c. The amount of premium to be amortized for the second semiannual interest payment period, using the interest method. Round to the nearest dollar. $

d. The amount of the bond interest expense for the first year. Round to the nearest dollar. $
Answer #1

a. Amount of cash proceeds from sale of bond = coupon interest x pvaf (i,n)+FV x pvif (i,n)

= We have, coupon interest = 50000000x0.12/2= $3000000

= I = 0.10/2 = 0.05

n = 5x2 = 10

FV = 50000000

Amount of cash proceeds = 3000000xpvaf(5%,8)+50000000xpvif(5%,8)

3000000x6.46321+50000000x0.67684 = 19389630+33842000 = $53231630

= Cash proceeds from sale of bind = $53231630

b. Amount of premium to be amortized for first semiannual interest payment period using interest method will be

Cash expense = 50000000x12%x6/12 = 3000000

Interest expense = 53231630x10%x6/12 = 2661582

Premium amortization= 3000000-2661582 = $338418

c. Current carrying value of bond = 53231630-338418 = $52893212

Cash interest = 50000000x12%x6/12 = 3000000

Interest expense = 52893212x10%x6/12 = 2644661

Amortization of premium for second semi annual interest period = 3000000-2644661 = $355339

d. The amount of bond interest expense for the first year = 2661582+2644661 = $5306243